1.The demand for some goods is set by formula Qd = 200-10Р, the supply is presented by the formula Qs = 60+15P 1) Compute the equilibrium price and quantity of goods at this market; 2) If a market price is fixed as Р = 10rub., calculate the demand and supply amount.


1
Expert's answer
2021-09-23T09:19:20-0400

"1) 200-10P = 60+15P"


"200-60 = 15P+10P"


"140 = 25P"


"P = \frac{140}{25}"


"P = 5.6"

Equilibrium price = 5.6



"Qd = 200-10(5.6)"


"= 200-56 = 144"


Equilibrium quantity = 144





2. When"P = 10rub"


demand amount = "200-10(10)"


"= 200-100 = 100"

demand amount = 100


Supply amount = "60+15(10)"


"60+150"


"= 210"


Supply amount = 210